OpenStudy (anonymous):

A disc jockey wants to select five songs from a new CD that contains 12 songs. How many ways can the disc jockey choose five songs

OpenStudy (anonymous):

12*5=60 So I believe 60 different ways (not 100% sure)

OpenStudy (kropot72):

The number of combinations of 12 different songs selected 5 at a time is \[12C5=\frac{12!}{5!7!}\]

OpenStudy (anonymous):

210?

OpenStudy (kropot72):

\[\frac{12!}{5!7!}=\frac{12\times11\times10\times9\times8}{5\times4\times3\times2\times1}=you\ can\ calculate\]

OpenStudy (anonymous):

yeah one sec

OpenStudy (anonymous):

95040/120

OpenStudy (anonymous):

Sorry for my wrong reply

OpenStudy (kropot72):

Correct so far. So what is the value of 95040/120 ?

OpenStudy (anonymous):

792

OpenStudy (kropot72):

Your answer is correct!

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)